Suppose a company has a stock price of $21.1 and has had earnings of $1.93 per share during the last twelve months. The consensus analyst forecast for earnings growth over the next five years is 8.9% per year. What is this stock's PEG, rounded to two decimal places?
Answer #1

The PEG ratio is computed by using the below formula:

= P/E Ratio / Earnings growth rate

P/E ratio is computed by using the below formula:

= MPS / EPS

= $ 21.1 / $ 1.93

= 10.93 Approximately

Earnings growth rate = 8.9%

By plugging these values in the above mentioned formula, we shall get:

= 10.93 / 8.9

= 1.23 Approximately

Feel free to ask in case of any query relating to this question